# A consistent, volume preserving, and adaptive mesh refinement-based framework for modeling non-isothermal gas-liquid-solid flows with phase change

### v1

@@ 3.1 Spatial discretization
The coarsest level of the grid covers the entire domain $\Omega$ with $N_{x}\times N_{y}$ rectangular cells. The cell size in the $x$ and $y$ directions is $\Delta x$ and $\Delta y$, respectively, as illustrated in Fig. 4(a). Unless stated otherwise, a uniform grid spacing $\Delta x=\Delta y=\Delta$ is used for all simulations in this work. We assume that the bottom-left corner of the computational domain $\Omega$ aligns with the origin $(0,0)$. The position of each center of the grid cell is given by $\bm{\mathrm{x}}_{i,j}=\left((i+\frac{1}{2})\Delta x,(j+\frac{1}{2})\Delta y\right)$, where $i=0,\ldots,N_{x}-1$ and $j=0,\ldots,N_{y}-1$. The face center in the $x-$direction, which is located half a grid space away from the cell center $\bm{\mathrm{x}}_{i,j}$ in the negative $x-$direction, is represented by $\bm{\mathrm{x}}_{i-\frac{1}{2},j}=\left(i\Delta x,(j+\frac{1}{2})\Delta y\right)$, where $i=0,\ldots,N_{x}$ and $j=0,\ldots,N_{y}-1$. Similar conventions apply to other face center locations. Scalar variables such as pressure ($p$), specific enthalpy ($h$), and temperature ($T$) are stored at the cell centers. The $x-$component of velocity ($u$) is stored at the centers of $x-$direction cell faces, while the $y-$component of velocity $v$ is stored at the centers of the faces of the $y-$ direction cell, as shown in Fig. 4(b). Material properties, including density ($\rho$), are stored at the cell centers along with interface tracking variables $\phi$ and $\varphi$. Second-order interpolation is used to interpolate the cell-centered quantities to face. Momentum forcing terms, such as surface tension, gravity, and Carman-Kozeny drag force, are stored at the cell faces. Standard second-order finite differences are employed to approximate spatial differential operators. The spatial discretizations of the key continuous operators are as follows:
@@
    The continuous strain rate tensor form of the viscous term is
    $\nabla\cdot\mbox{}\left[\mu\left(\nabla\bm{\mathrm{u}}+\nabla\bm{\mathrm{u}}^{\intercal}\right)\right]=\left[\begin{array}[]{c}2\frac{\partial}{\partial x}\left(\mu\frac{\partial u}{\partial x}\right)+\frac{\partial}{\partial y}\left(\mu\frac{\partial u}{\partial y}+\mu\frac{\partial v}{\partial x}\right)\\ 2\frac{\partial}{\partial y}\left(\mu\frac{\partial v}{\partial y}\right)+\frac{\partial}{\partial x}\left(\mu\frac{\partial v}{\partial x}+\mu\frac{\partial u}{\partial y}\right)\\ \end{array}\right],$ (29)
    which couples the velocity components through spatially variable viscosity
    $\bm{\mathrm{L_{\mu}}}\bm{\mathrm{u}}=\left[\begin{array}[]{c}(\bm{\mathrm{L_{\mu}}}\bm{\mathrm{u}})^{x}_{i-\frac{1}{2},j}\\ (\bm{\mathrm{L_{\mu}}}\bm{\mathrm{u}})^{y}_{i,j-\frac{1}{2}}\\ \end{array}\right].$ (30)
@@ 3.2 Temporal discretization
In this section, we detail the temporal discretization method utilized in our study for the continuous equations of motion. Within each time step [$t^{n},t^{n+1}$], we employ $n_{\text{cycles}}$ cycles of fixed-point iteration to approximate the solution to the coupled fluid-thermal system. At the start of the simulation ($t=0$), all variables are initialized with the initial conditions of the problem. At the beginning of the each time step, with cycle number $k=0$, variables are initialized to their values from the previous time step, denoted as $\bm{\mathrm{u}}^{n+1,0}=\bm{\mathrm{u}}^{n}$, $p^{n+\frac{1}{2},0}=p^{n-\frac{1}{2}}$, $h^{n+1,0}=h^{n}$, $\phi^{n+1,0}=\phi^{n}$, and $\varphi^{n+1,0}=\varphi^{n}$, and iterated till $k=n_{\text{cycles}}-1$. Convective and surface tensions terms are treated explicitly, while other terms are treated implicitly. Unless otherwise specified, $n_{\text{cycles}}=2$ is used in this work.
@@ 3.2.1 Consistent mass-momentum integrators
For incompressible multiphase flows, mass is advected indirectly via an interface tracking variable. In the present work this corresponds to the advection of the level set variable $\phi$. The momentum equation also contains a mass flux term $\bm{\mathrm{m_{\rho}}}\equiv\rho\bm{\mathrm{u}}$ in the the convective operator $\nabla\cdot\mbox{}(\rho\bm{\mathrm{u}}\otimes\bm{\mathrm{u}})=\nabla\cdot\mbox{}(\bm{\mathrm{m_{\rho}}}\otimes\bm{\mathrm{u}})$. For high density ratio flows $\rho^{\rm L}/\rho^{\rm G}>100$ this weak coupling of mass and momentum leads to numerical instabilities. One way to ensure strong coupling between mass and momentum advection is to solve the redundant mass balance Eq. (15a) and use the same mass flux $\bm{\mathrm{m_{\rho}}}$ in the two advective operators: $\nabla\cdot\mbox{}(\rho\bm{\mathrm{u}})$ and $\nabla\cdot\mbox{}(\rho\bm{\mathrm{u}}\otimes\bm{\mathrm{u}})$. This idea was proposed in our prior works [29, 30].
@@
and $\breve{\bm{\mathrm{\rho}}}^{n+1,k+1}$ is the updated value of the density obtained from the last stage of the SSP-RK3 integrator. The term $A_{d}=C_{d}\frac{\displaystyle\varphi_{\text{S}}^{2}}{\displaystyle(1-\varphi_{\text{S}})^{3}}+\epsilon$ in the momentum equation represents the Carman-Kozeny drag coefficient that is used to retard any flow in the solid domain. Here, $\varphi_{\text{S}}=H(1-\varphi)$ is the solid fraction of the grid cell, and $\epsilon=10^{-3}$ is a small number to avoid a division by zero and to control the strength of penalty parameter ($C_{d}/\epsilon$) in the solid region. In the solid domain, $C_{d}$ is also taken to be large in magnitude to retard fluid motion. When equating drag force with the first term on the left hand side of the momentum equation, a sufficiently large value is obtained for $C_{d}=\rho^{\rm S}/\Delta t$. The three-phase surface tension force is evaluated using the midpoint values of the level set function $\phi^{n+\frac{1}{2},k+1}=\frac{1}{2}(\phi^{n+1,k+1}+\phi^{n})$.
@@
A test problem to check consistency of mass-momentum transport: To illustrate the importance of consistent mass and momentum transport for high density ratio flows, consider an advection of a dense, isothermal bubble moving in a uniform velocity $\bm{\mathrm{u}}_{c}=(u,v)=(1,1)$ in a square periodic domain $\Omega\in[0,1]^{2}$. The radius of the bubble is $R=0.2$ and it is initially centered at the middle of the domain: $(x_{i}=0.5,y_{i}=0.5)$. The density ratio between the bubble and the outer liquid is $\rho_{i}/\rho_{o}=10,000$. The domain is discretized with 2 levels of mesh, with the coarsest mesh size of $N_{x}\times N_{y}=128^{2}$. A refinement ratio of $n_{\rm ref}=2$ is used between the two levels. The bubble’s interface is always placed on the finest level. Further details on the adaptive mesh refinement (AMR) are provided in the next section. Viscosity $\mu$ is set to zero in the entire domain, and no other body force (e.g., gravity or surface tension) is present. Under these conditions, the governing equations for this problem reduce to
$\frac{\partial\phi}{\partial t}+\bm{\mathrm{u}}\cdot\nabla{\phi}=0,$ (41a)
@@
For this problem it is expected that the bubble advects with a constant velocity $\bm{\mathrm{u}}_{c}$ and returns to its original position (with minimal distortions) after $\tau=d/|\bm{\mathrm{u}}_{c}|$, in which $d$ represents the distance travelled to reach its original position. We simulate the advection of the bubble considering four scenarios:
@@
in which $\beta=\int_{\Omega}\rho\,\rm{d}V$ and $\mathcal{B}=\mathcal{M}$ denote the relative change in the mass of the system, and $\beta=\int_{\Omega}\rho\bm{\mathrm{u}}\,\rm{d}V$ and $\mathcal{B}=\mathcal{L}_{\bm{\mathrm{x}}}$ denote the relative change in the linear momentum of the system. For this problem we expect that the mass and momentum of the system should remain invariant over time. In order to check whether any catastrophic instability occurs in the four cases, we run the simulations till $t=10$, which corresponds to 10 complete passes of the dense bubble across the domain.
@@
Fig. 5 shows the results for the four cases. It can be observed that Case A produces spurious changes in mass and momentum of the system that increase slowly over time. However, adding the residual force in the momentum equation (Case B) reduces the spurious changes in mass and momentum, and keeps them relatively low and constant over time. Similarly, using the same time integrator (Case C and Case D) also yields very small percentage changes in both mass and momentum of the system that remain steady over time. In all four cases, the simulation remained stable till $t=10$. The dynamics of the problem are illustrated in Fig. 8 of the results Sec. 7. Based on the results of this test, we choose the integrator combination of Case D (i.e., RK-2 integrator for integrating both mass and momentum equations and considering an additional residual force in the momentum equation) in this work.
@@ 4 Adaptive mesh refinement (AMR)
    Tagging cells based on the signed distance function: To resolve the gas-PCM interface with sufficient mesh resolution, a value-based tagging criterion is used. A cell is tagged for grid refinement when its signed distance function value is within zero threshold, e.g., $|\phi|\leq 2\Delta x_{0}$. Using numerical simulations, we find that this amount of tagging is sufficient to resolve the sharp variation in material properties ($\rho$, $\kappa$, $\mu$) across the gas-PCM interface.
@@
The Stefan problem is simulated by considering a quasi one-dimensional computational domain $\Omega\in[0,1]\times[0,0.05]$ with $l=2$ mesh levels and a refinement ratio of $n_{\text{ref}}=2$. The coarsest level ($l=1$) is covered by $N_{x0}\times N_{y0}=640\times 32$ number of cells. The simulation runs until $t=10$ s with a constant time step size of $\Delta t=10^{-4}$ s. The domain is periodic in the $y$-direction. Initially, the liquid PCM occupies the entire domain and has a temperature of $T_{i}=973.6$ K. The temperature at the left boundary ($x=0$) is set at $T_{o}=298.6$ K to induce solidification. The melting/fusion temperature of the PCM is $T_{m}=933.6$ K. The right end of the boundary ($x=1$) is kept adiabatic (homogeneous Neumann boundary condition for temperature). The flow solver uses zero-velocity and zero-pressure/outflow boundary conditions at the left and right ends, respectively. The phase change material’s thermophysical properties, largely aluminum-based, are listed in Table 1. Here, the liquid and solid densities are assumed to be $\rho^{\rm L}=2700$ and $\rho^{\rm S}=500$ kg/m3, respectively, and both fluid and solid viscosities are set to zero. Upon solidification, the material expands due to its lower density than the liquid phase. This problem has been solved both analytically and numerically (on uniform grids) in our previous work [1]. We use this problem to analyze the performance of the two tagging methods ($\varphi$- vs. $\nabla\varphi$-based) to identify the mushy region effectively within the AMR framework.
@@ 5 Complete solution algorithm
    First, we advect the level set function $\phi^{n}$ using the (non-divergence-free) velocity field $\mathbf{u}$ to obtain $\phi^{n+1,k+1}$:
    $\frac{\phi^{n+1,k+1}-\phi^{n}}{\Delta t}+\left(\nabla\cdot(\phi\mathbf{u})\right)^{n+1,k}=(\phi\nabla\cdot\mathbf{u})^{n+1,k}.$ (49)
@@
    Here, $H^{n}$ is computed from the level set function $\phi^{n}$ using Eq. (4). We remark that explicitly advecting $H$ is not essential; instead directly computing $H^{n+1,k+1}$ from $\phi^{n+1,k+1}$ (through Eq. (4)) is sufficient. However, as demonstrated above for mass flux, the advective flux of Heaviside variable $H\bm{\mathrm{u}}$ could be used to advect additional scalar variables of a more involved problem.
@@ 6 Software implementation
The low Mach enthalpy method and its proposed improvements described in this work have been implemented within the IBAMR library [36], an open-source C++ software enabling the simulation of CFD and fluid-structure interaction problems. IBAMR utilizes adaptive mesh refinement (AMR) and Cartesian grid management through the SAMRAI framework [37, 38], with linear solver support provided by the PETSc library ([39, 40]). All simulations presented in this work are implemented in the following GitHub repository:https://github.com/IBAMR/IBAMR/tree/ramakrishnan/phase_change_amr_mass_fix.

### v2

@@ 3.1 Spatial discretization
The coarsest level of the grid covers the entire domain $\Omega$ with $N_{x}\times N_{y}$ rectangular cells. The cell size in the $x$ and $y$ directions is $\Delta x$ and $\Delta y$, respectively, as illustrated in Fig. 4(a). Unless stated otherwise, a uniform grid spacing $\Delta x=\Delta y=\Delta$ is used for all simulations in this work. We assume that the bottom-left corner of the computational domain $\Omega$ aligns with the origin $(0,0)$. The position of each center of the grid cell is given by $\bm{\mathrm{x}}_{i,j}=\left((i+\frac{1}{2})\Delta x,(j+\frac{1}{2})\Delta y\right)$, where $i=0,\ldots,N_{x}-1$ and $j=0,\ldots,N_{y}-1$. The face center in the $x-$direction, which is located half a grid space away from the cell center $\bm{\mathrm{x}}_{i,j}$ in the negative $x-$direction, is represented by $\bm{\mathrm{x}}_{i-\frac{1}{2},j}=\left(i\Delta x,(j+\frac{1}{2})\Delta y\right)$, where $i=0,\ldots,N_{x}$ and $j=0,\ldots,N_{y}-1$. Similar conventions apply to other face center locations. Scalar variables such as pressure ($p$), specific enthalpy ($h$), and temperature ($T$) are stored at the cell centers. The $x-$component of velocity ($u$) is stored at the centers of $x-$direction cell faces, while the $y-$component of velocity $v$ is stored at the centers of the faces of the $y-$ direction cell, as shown in Fig. 4(b). Material properties, including density ($\rho$), are stored at the cell centers along with interface tracking variables ${\color[rgb]{0,0,0}\Phi}$ and $\varphi$. Second-order interpolation is used to interpolate the cell-centered quantities to face. Momentum forcing terms, such as surface tension, gravity, and Carman-Kozeny drag force, are stored at the cell faces. Standard second-order finite differences are employed to approximate spatial differential operators. The spatial discretizations of the key continuous operators are as follows:
@@
    The continuous strain rate tensor form of the viscous term, which couples the velocity components through spatially variable viscosity, is given by
    $\nabla\cdot\mbox{}\left[\mu\left(\nabla\bm{\mathrm{u}}+\nabla\bm{\mathrm{u}}^{\intercal}\right)\right]=\left[\begin{array}[]{c}(\bm{\mathrm{L_{\mu}}}\bm{\mathrm{u}})^{x}_{i-\frac{1}{2},j}\\ (\bm{\mathrm{L_{\mu}}}\bm{\mathrm{u}})^{y}_{i,j-\frac{1}{2}}\\ \end{array}\right]=\left[\begin{array}[]{c}2\frac{\partial}{\partial x}\left(\mu\frac{\partial u}{\partial x}\right)+\frac{\partial}{\partial y}\left(\mu\frac{\partial u}{\partial y}+\mu\frac{\partial v}{\partial x}\right)\\ 2\frac{\partial}{\partial y}\left(\mu\frac{\partial v}{\partial y}\right)+\frac{\partial}{\partial x}\left(\mu\frac{\partial v}{\partial x}+\mu\frac{\partial u}{\partial y}\right)\\ \end{array}\right].$ (31)
@@ 3.2 Temporal discretization
In this section, we detail the temporal discretization method utilized in our study for the continuous equations of motion. Within each time step [$t^{n},t^{n+1}$], we employ $n_{\text{cycles}}$ cycles of fixed-point iteration to approximate the solution to the coupled fluid-thermal system. At the start of the simulation ($t=0$), all variables are initialized with the initial conditions of the problem. At the beginning of the each time step, with cycle number $k=0$, variables are initialized to their values from the previous time step, denoted as $\bm{\mathrm{u}}^{n+1,0}=\bm{\mathrm{u}}^{n}$, $p^{n+\frac{1}{2},0}=p^{n-\frac{1}{2}}$, $h^{n+1,0}=h^{n}$, ${\color[rgb]{0,0,0}\Phi}^{n+1,0}={\color[rgb]{0,0,0}\Phi}^{n}$, and $\varphi^{n+1,0}=\varphi^{n}$, and iterated till $k=n_{\text{cycles}}-1$. Convective and surface tensions terms are treated explicitly, while other terms are treated implicitly. Unless otherwise specified, $n_{\text{cycles}}=2$ is used in this work.
@@ 3.2.1 Consistent mass-momentum integrators
For incompressible multiphase flows, mass is advected indirectly via an interface tracking variable. In the present work this corresponds to the advection of the level set variable ${\color[rgb]{0,0,0}\Phi}$. The momentum equation also contains a mass flux term $\bm{\mathrm{m_{\rho}}}\equiv\rho\bm{\mathrm{u}}$ in the the convective operator $\nabla\cdot\mbox{}(\rho\bm{\mathrm{u}}\otimes\bm{\mathrm{u}})=\nabla\cdot\mbox{}(\bm{\mathrm{m_{\rho}}}\otimes\bm{\mathrm{u}})$. For high density ratio flows $\rho^{\rm L}/\rho^{\rm G}>100$ this weak coupling of mass and momentum leads to numerical instabilities. One way to ensure strong coupling between mass and momentum advection is to solve the redundant mass balance Eq. (16a) and use the same mass flux $\bm{\mathrm{m_{\rho}}}$ in the two advective operators: $\nabla\cdot\mbox{}(\rho\bm{\mathrm{u}})$ and $\nabla\cdot\mbox{}(\rho\bm{\mathrm{u}}\otimes\bm{\mathrm{u}})$. This idea was proposed by Desjardins and Moureau [37] in 2010 and later extended to achieve second-order accuracy in our prior works [29,30].
@@
and $\breve{\bm{\mathrm{\rho}}}^{n+1,k+1}$ is the updated value of the density obtained from the last stage of the SSP-RK3 integrator. The term $A_{d}=C_{d}\frac{\displaystyle\varphi_{\text{S}}^{2}}{\displaystyle(1-\varphi_{\text{S}})^{3}+\epsilon}$ in the momentum equation represents the Carman-Kozeny drag coefficient that is used to retard any flow in the solid domain. Here, $\varphi_{\text{S}}=H(1-\varphi)$ is the solid fraction of the grid cell, and $\epsilon=10^{-3}$ is a small number to avoid a division by zero and to control the strength of penalty parameter ($C_{d}/\epsilon$) in the solid region. In the solid domain, $C_{d}$ is also taken to be large in magnitude to retard fluid motion. When equating drag force with the first term on the left hand side of the momentum equation, a sufficiently large value is obtained for $C_{d}=\rho^{\rm S}/\Delta t$. The three-phase surface tension force is evaluated using the midpoint values of the level set function ${\color[rgb]{0,0,0}\Phi}^{n+\frac{1}{2},k+1}=\frac{1}{2}({\color[rgb]{0,0,0}\Phi}^{n+1,k+1}+{\color[rgb]{0,0,0}\Phi}^{n})$.
@@
A test problem to check consistency of mass-momentum transport: To illustrate the importance of consistent mass and momentum transport for high density ratio flows, consider an advection of a dense, isothermal droplet [^1] This benchmark test is provided in IBAMR GitHub within the directory examples/phase_change/ex2. moving in a uniform velocity $\bm{\mathrm{u}}_{c}=(u,v)=(1,1)$ in a square periodic domain $\Omega\in[0,1]^{2}$. The radius of the droplet is $R=0.2$ and it is initially centered at the middle of the domain: $(x_{i}=0.5,y_{i}=0.5)$. The density ratio between the droplet and the outer liquid is $\rho_{i}/\rho_{o}=10,000$. The domain is discretized with 2 levels of mesh, with the coarsest mesh size of $N_{x}\times N_{y}=128^{2}$. A refinement ratio of $n_{\rm ref}=2$ is used between the two levels. The droplet’s interface is always placed on the finest level. Further details on the adaptive mesh refinement (AMR) are provided in the next section. Viscosity $\mu$ is set to zero in the entire domain, and no other body force (e.g., gravity or surface tension) is present. Under these conditions, the governing equations for this problem reduce to
$\frac{\partial{\color[rgb]{0,0,0}\Phi}}{\partial t}+\bm{\mathrm{u}}\cdot\nabla{{\color[rgb]{0,0,0}\Phi}}=0,$ (42a)
@@
For this problem it is expected that the droplet advects with a constant velocity $\bm{\mathrm{u}}_{c}$ and returns to its original position (with minimal distortions) after $\tau=d/|\bm{\mathrm{u}}_{c}|$, in which $d$ represents the distance travelled to reach its original position. We simulate the advection of the droplet considering four scenarios:
@@
in which $\beta=\int_{\Omega}\rho\,\rm{d}V$ and $\mathcal{B}=\mathcal{M}$ denote the relative change in the mass of the system, and $\beta=\int_{\Omega}\rho\bm{\mathrm{u}}\,\rm{d}V$ and $\mathcal{B}=\mathcal{L}_{\bm{\mathrm{x}}}$ denote the relative change in the linear momentum of the system. $\beta_{0}$ represents quantities computed using the system’s initial conditions. For this problem we expect that the mass and momentum of the system should remain invariant over time. In order to check whether any catastrophic instability occurs in the four cases, we run the simulations till $t=10$, which corresponds to 10 complete passes of the droplet across the domain.
@@
Fig. 5 shows the results for the four cases. It can be observed that Case A produces spurious changes in mass and momentum of the system that increase slowly over time. However, adding the residual force in the momentum equation (Case B) reduces the spurious changes in mass and momentum, and keeps them relatively low and constant over time. Similarly, using the same time integrator (Case C and Case D) also yields very small percentage changes in both mass and momentum of the system that remain steady over time. The spurious changes in mass and momentum for Cases B, C, and D have similar magnitudes, as shown in Fig. 6. In all four cases, the simulation remained stable till $t=10$. The dynamics of the problem are illustrated in Fig. 9 of the results Sec. 7. Based on the results of this test, we choose the integrator combination of Case D (i.e., RK-2 integrator for integrating both mass and momentum equations and considering an additional residual force in the momentum equation) in this work.
Figure 6: Percentage change in (A) mass (B) x-momentum and (C) the y-momentum of the system for various choices of time integrators: Case B - RK-2 integrator for momentum equation, SSP-RK3 integrator for mass balance equation, and residual force $\mathcal{R}\bm{\mathrm{u}}$ in the momentum equation; Case C - RK-2 integrator for both mass and momentum equations; Case D - RK-2 integrator for both mass and momentum equations, and residual force $\mathcal{R}\bm{\mathrm{u}}$ in the momentum equation. $n_{\text{cycles}}=2$ is used for all four cases.
@@ 4 Adaptive mesh refinement (AMR)
    Tagging cells based on the signed distance function: To resolve the gas-PCM interface with sufficient mesh resolution, a value-based tagging criterion is used. A cell is tagged for grid refinement when its signed distance function value is within zero threshold, e.g., $|{\color[rgb]{0,0,0}\Phi}|\leq 2\Delta x_{0}$. Using numerical simulations, we find that this amount of tagging is sufficient to resolve the sharp variation in material properties ($\rho$, $\kappa$, $\mu$) across the gas-PCM interface.
@@
[…]
@@ 5 Complete solution algorithm
    First, we advect the level set function ${\color[rgb]{0,0,0}\Phi}^{n}$ using the (non-divergence-free) velocity field $\mathbf{u}$ to obtain ${\color[rgb]{0,0,0}\Phi}^{n+1,k+1}$:
    $\frac{{\color[rgb]{0,0,0}\Phi}^{n+1,k+1}-{\color[rgb]{0,0,0}\Phi}^{n}}{\Delta t}+\left(\nabla\cdot({\color[rgb]{0,0,0}\Phi}\mathbf{u})\right)^{n+1,k}=({\color[rgb]{0,0,0}\Phi}\nabla\cdot\mathbf{u})^{n+1,k}.$ (50)
@@
    Here, $H^{n}$ is computed from the level set function ${\color[rgb]{0,0,0}\Phi}^{n}$ using Eq. (5). We remark that explicitly advecting $H$ is not essential; instead directly computing $H^{n+1,k+1}$ from ${\color[rgb]{0,0,0}\Phi}^{n+1,k+1}$ (through Eq. (5)) is sufficient. However, as demonstrated above for mass flux, the advective flux of Heaviside variable $H\bm{\mathrm{u}}$ could be used to advect additional scalar variables of a more involved problem.
@@ 6 Software implementation
The low Mach enthalpy method and its proposed improvements described in this work have been implemented within the IBAMR library [41], an open-source C++ software enabling the simulation of CFD and fluid-structure interaction problems. IBAMR utilizes adaptive mesh refinement (AMR) and Cartesian grid management through the SAMRAI framework [42, 43], with linear solver support provided by the PETSc library ([44, 45]). All simulations presented in this work are implemented in the following GitHub repository:https://github.com/IBAMR/IBAMR/tree/phase_change_patch_6.